An individual has 40,000 in income per year. The person will get sick with probability 0.1. If he does get sick, the medical bills will total 30,000. The following tables shows the utility derived from certain amounts of income:
Income Utility
40,000 200
37,000 195
35,000 190
30,000 170
20,000 140
10,000 100
Calculate the actuarially fair premium.
1
Expert's answer
2019-11-11T15:31:33-0500

The actuarially fair premium is:

Paf = Loss×P(loss) = 30,000×0.1 = 3,000.